Boston Railroad decided to use the high-low method and operating data from the past six months to estimate the fixed and variable components of transportation costs. The activity base used by Boston Railroad is a measure of railroad operating activity, termed "gross-ton miles," which is the total number of tons multiplied by the miles moved.

Transportation Costs Gross-Ton Miles
January $669,600 243,000
February 746,500 272,000
March 527,600 176,000
April 715,700 263,000
May 600,300 212,000
June 769,600 286,000

Determine the variable cost per gross-ton mile and the total fixed cost.

Variable cost (Round to two decimal places.) $ per gross-ton mile
Total fixed cost $
Answer #1

Solution:

Variable cost per gross ton mile = (Cost at high level - Cost at low level) / (High level activity - Low level activity)

= (($769,600 - $527,600) / (286,000 - 176,000) = $242,000 / 110000 = $2.20 per gross ton mile

Total fixed costs = $769,600 - (286000*$2.20) = $140,400